OpenStudy (anonymous):

Express the 4th-order differential equation as a system of first order differential equations using matrix notation

OpenStudy (anonymous):

\[y^{(4)} +\frac{ y^{(3)} }{t^2 }+\sin(t)y''+e^ty=t^2\]

OpenStudy (anonymous):

i understand that i have to substitute the y derivatives as other variables, but i am really confused on how to convert it to a matrix

OpenStudy (anonymous):

so far i have: \[x_1=y, x_2=y', x_3=y'', x_4 = y^{(3)}\]

OpenStudy (anonymous):

which gives me\[x_4'+\frac{ x_4 }{t^2 }+\sin(t)x_3 +e^tx_1 = t^2\] so how do convert this using matrix notation?

OpenStudy (anonymous):

Using the substitution you've made, you also have that \[\begin{cases} x_1=y\\ x_2=x_1'=y'\\ x_3=x_2'=y''\\ x_4=x_3'=y^{(3)} \end{cases}\] So you have the system of equations \[\large\begin{cases} x_1'=x_2\\\\ x_2'=x_3\\\\ x_3'=x_4\\\\ x_4'+\frac{1}{t^2}x_4+\sin t~x_3+e^t~x_1=t^2 \end{cases}\]

OpenStudy (anonymous):

It shouldn't be too difficult to write that in matrix form.

OpenStudy (anonymous):

okay thank you, i was confused that you had to make all the terms as derivative equations since your x'_4 is

OpenStudy (anonymous):

You're welcome.